A new car is purchased for 15500 dollars. The value of the car depreciates at 7.5% per year. What will the value of the car be, to the nearest cent, after 6 years?



1
Expert's answer
2021-05-11T09:01:28-0400

"V=P(1-r)^x"

V is the depreciated value

P is the original value P=15500$

r is the rate of depreciation in decimal r=7.5%=0.075

x is Number of Time Periods x=6 years

"V=15500(1-0.075)^6=" 9709.17$